In Chemistry, you mixed a 10% solution with a 15% solution to get 100 milliliters of a 12% solution. How much of the 10% solution did you you use in the mixture?
Answer by TimothyLamb(4379)

---
10x + 15y = 12*100
x + y = 100
---
10x + 15y = 1200
x + y = 100
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
x= 60 ml of 10%
y= 40 ml of 15%
---
answer:
60 ml of 10%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php